OpenStudy (anonymous):

use the binomial theoroem to find the coefficient of x^7 in the expansion of (x+1)^11

OpenStudy (raden):

(11-4)C4 = 7C4

OpenStudy (anonymous):

still confused

OpenStudy (raden):

use the formula of combination nCr = n!/{(n-r)!r!}

OpenStudy (anonymous):

11*10*9*8/7*6*5*4*3*2*1?

OpenStudy (raden):

7C3, it means n=7 and r=3 now apply of them to the formula above

OpenStudy (anonymous):

ok

OpenStudy (raden):

opss.. 7C4 i meant :) so, n=7 and r=4

OpenStudy (anonymous):

still confused

OpenStudy (raden):

did u know about factorial

OpenStudy (anonymous):

a little

OpenStudy (raden):

what's mean of 7! ?

OpenStudy (anonymous):

you times 7 all the way down to 1 or whatever the other one is

OpenStudy (raden):

ok, so what's the result of 7!/{(7-4)!4!}

OpenStudy (raden):

it similar by 7!/(3!4!)

OpenStudy (anonymous):

7*6*5, 4*3*2*1

OpenStudy (raden):

well, i hope u can undertanding this... 7! / 3!4! = (7*6*5*4!)! / (3*2*1*4!) (cancel out for 4!) = 7*6*5 / 3*2*1 = 210/6=35

OpenStudy (anonymous):

so is 35 my answer

OpenStudy (raden):

yes

OpenStudy (anonymous):

not right im suppose to have something than x^7

OpenStudy (raden):

u have asked to find the coefficient of x^7, so the answer is 35 completely 35x^7 :P

OpenStudy (anonymous):

it says thats not right

OpenStudy (raden):

im very sure 1000000000000000^infinity %, my answer was right :P except, u have typo in ur question

OpenStudy (anonymous):

no my question is right

OpenStudy (raden):

do u have choices

OpenStudy (anonymous):

no

OpenStudy (raden):

ahhhhhhaaaa.... damned, i have mistaken look that numbers it should be, x^(11-4)*(1)4 = x^7 so, the coefficient of x^7 is ONE sorry, man :o

OpenStudy (anonymous):

ok so what whats the answer

OpenStudy (raden):

wait... i think i was do mistake again

OpenStudy (raden):

it should be 11C4 not 7C4 so, 11C4 = 11!/(11-4)!4! = 11!/8!4! = 11*10*9*8!/(8!*4*3*2*1) = 330 the answer is 330

OpenStudy (raden):

is it clear